Circular No. 6683

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970616 V. Connaughton, National Academy of Sciences/National Research

Council and Marshall Space Flight Center (MSFC); R. M. Kippen, R. Preece, and G. N. Pendleton, University of Alabama at Huntsville and MSFC; and S. D. Barthelmy, Universities Space Research Association (USRA) and Goddard Space Flight Center (GSFC), report for the BATSE team: “The transient x-ray XTE source reported below by Marshall et al. lies in the error box of a γ-ray burst seen by BATSE on June 16.757 UT. The Rapid Burst Reponse location, which was distributed through the GRB Coordinates Network/BATSE Coordinates Distribution Network, 20 min after the initial trigger, was α = 1^{h}21^{m}51^{s}, δ = −7^{o}09′.4 (equinox 2000.0), with a 1σ error radius of 0^{o}.3 statistical and ∼ 2^{o} systematic. The burst was multipeaked, with a total duration of ∼ 200 s. A maximum flux of 23.8 ± 0.5 photons cm^{−2} s^{−1} (above 20 keV) was measured 90 s after the trigger time, placing this event in the brightest 2 percent of all BATSE bursts. Its total fluence above 20 keV is estimated to be 4.01 ± 0.14 × 10^{−5} erg cm^{−2}.”

F. E. Marshall, GSFC; T. Takeshima, and S. D. Barthelmy, USRA and GSFC; C. R. Robinson, USRA and MSFC; and K. Hurley, University of California at Berkeley, report the detection of an x-ray afterglow for the γray burst detected with BATSE on June 16.7568 UT. Scanning observations with the Proportional Counter Array on the Rossi X-ray Timing Explorer (RXTE) 4.0 hr after the burst revealed a source at α = 1^{h}18^{m}.9 ± 0^{m}.7 (90-percent statistical confidence), δ = −5^{o}30′ ± 18′ (equinox 2000.0) with a countrate of 4.5 ± 1.6 counts/s or 1.1 × 10^{−11} erg s^{−1} cm^{−2} in the band 2–10 keV. It is the brightest source detected within the 2^{o} BATSE error circle, and it also overlaps the preliminary inter-planetary network (IPN) annulus determined from the time difference in the BATSE and Ulysses detections. The IPN annulus has a radius of 49^{o}.808, is centered at α = 22^{h}10^{m}30^{s}, δ = −27^{o}24′.8, and has a full width of 4′.6. An annulus with a smaller width will be available after further processing. The intersection of the error regions has corners at 1^{h}19^{m}.6, −5^{o}49′; 1^{h}19^{m}.6, −5^{o}40′; 1^{h}18^{m}.4, −5^{o}12′; 1^{h}18^{m}.2, −5^{o}16′. A subsequent RXTE observation on June 18.08 detected no significant flux from this position, with a preliminary 90-percent confidence upper limit of 2.5 counts/s.

1997 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green